# Multi-Preference Lambda-weighted Listwise DPO for Small-Scale Model Alignment

### v5

@@ Introduction
$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}|x)-P_{\theta}(y_{i}|x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)\right]$ (11)
This gradient updates the model to increase likelihood of responses aligned with weighted human preferences, while decreasing likelihood of responses overrepresented in the reference policy. Importantly, when $\lambda$ is fixed (e.g., $\lambda=(1,0,0)$), the objective reduces to a single-objective listwise DPO loss.
Lambda-DPO Pipeline.
@@
In practice, we can reuse datasets like UltraFeedback (Du et al. 2023), where multiple preference criteria are annotated. During inference, changing $\lambda$ enables personalized alignment without modifying model parameters.
Theoretical Analysis of Multi-Preference Lambda-weighted Listwise DPO
@@
While Direct Preference Optimization (DPO) offers a principled and scalable approach to aligning language models with human feedback, it is fundamentally limited by its reliance on pairwise supervision. In this section, we provide a detailed theoretical analysis of these limitations and motivate our proposed extensions—namely, listwise supervision and preference mixing via simplex weighting—to overcome them in multi-objective alignment settings.
Pairwise DPO and Its Limitations.
@@
Given a prompt $x$ and a pair of sampled completions $y_{w}$ (preferred) and $y_{l}$ (dispreferred), DPO defines a binary target distribution over the pair as:
$p^{*}(y_{i}\mid x)=\begin{cases}1&\text{if }y_{i}=y_{w}\\ 0&\text{if }y_{i}=y_{l}\end{cases}$ (12)
This construction is simple and efficient, allowing the loss to be computed from binary human preferences. However, it treats the preference signal as a hard decision, discarding nuanced feedback such as “somewhat better” or “equally good.” Furthermore, the normalization over only two samples induces high-variance gradients, as the updates depend on stochastic comparisons between limited completions. Since only the relative preference between $y_{w}$ and $y_{l}$ is modeled, all other completions (which might be equally or more suitable) are ignored. This can lead to suboptimal generalization, especially when the number of plausible high-quality completions is large.
Listwise Distribution and Multi-Preference Supervision.
To address the expressiveness gap, we generalize DPO to operate over a list of completions $\{y_{1},\dots,y_{N}\}$ using listwise supervision. Specifically, given a reward model $r(x,y_{i})$ that quantifies the quality of each completion, we define a normalized target distribution over the list using softmax:
$p^{*}(y_{i}\mid x)=\frac{\exp(r(x,y_{i}))}{\sum_{j=1}^{N}\exp(r(x,y_{j}))}$ (13)
This formulation assigns probability mass proportionally to reward values, providing smoother gradients and improving credit assignment. Unlike the binary setting, listwise DPO makes use of the full ranking structure, offering a richer training signal and faster convergence.
Moreover, human preferences in practical applications are inherently multi-dimensional—comprising objectives like helpfulness, safety, factuality, politeness, and conciseness. Suppose we define $m$ distinct reward models, each representing one objective. We can construct a family of listwise distributions $\{p^{*(k)}(y_{i}|x)\}_{k=1}^{m}$, each aligned with a single criterion. A convex combination of these objectives is parameterized by a preference vector $\lambda\in\Delta^{m}$, forming a new composite distribution:
$p^{\lambda}(y_{i}\mid x)=\sum_{k=1}^{m}\lambda_{k}\cdot p^{*(k)}(y_{i}\mid x)$ (14)
This formulation enables controllable alignment by adjusting $\lambda$, allowing practitioners to emphasize certain values over others (e.g., prioritize safety over informativeness). The model is trained to match this target distribution via cross-entropy:
$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\log P_{\theta}(y_{i}\mid x)\right]$ (15)
Its gradient reveals the direction of improvement, given by $\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=$
$-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\right]$ (16)
which encourages the model’s policy distribution to match the aggregated preference signal from multiple human value dimensions.
Summary.
In summary, Multi-Preference Lambda-weighted Listwise DPO offers a theoretically grounded generalization of the original DPO framework. First, it extends binary supervision to listwise distributions, providing smoother and more informative gradient updates. Second, it integrates diverse human objectives through a simplex-weighted mixture of reward-aligned targets, enabling dynamic and tunable preference alignment. Finally, it reduces sample inefficiency and training instability by leveraging full output rankings and avoiding overfitting to binary labels, thus making it well-suited for real-world alignment tasks with conflicting value axes.
Comparative Experiment
We compare standard DPO and our proposed $\lambda$-DPO under identical training and evaluation setups on two models: LLaMA 3.2-1B Instruct and Qwen 1.8B Chat. Both models are first supervised fine-tuned (SFT) on UltraChat-200k, then aligned using the UltraFeedback listwise dataset.
Training Details.
We use LoRA (rank 8, all-layer target) with a learning rate of $5\times 10^{-6}$, batch size 1, gradient accumulation steps 8, cosine learning rate schedule, and 5 epochs. The KL penalty $\beta$ is set to 0.1. $\lambda$-DPO uses the lambda_dpo loss with chunk size 1.
Main Results
Table 1 reports the accuracy of DPO and $\lambda$-DPO on six representative benchmarks for both LLaMA 3.2-1B Instruct and Qwen 1.8B Chat. On LLaMA, $\lambda$-DPO consistently improves performance across all tasks, with notable gains such as +1.1% on ARC, +3.3% on GSM8K, and +3.6% on TruthfulQA. The improvement on WSC273 is also evident (+1.8%). For Qwen 1.8B Chat, $\lambda$-DPO outperforms DPO on all benchmarks except TruthfulQA, where DPO achieves a slightly higher score. Notably, large gains are observed on GSM8K (+14.7%) and MMLU (+3.9%), demonstrating the generalization benefits of $\lambda$-DPO across different model families.
Model Method MMLU ARC HellaSwag TruthfulQA WSC273 GSM8K
LLaMA 3.2-1B Instruct ORPO 0.456 0.355 0.447 0.435 0.678 0.327
SIMPO 0.459 0.362 0.455 0.381 0.703 0.318
DPO 0.458 0.364 0.456 0.378 0.696 0.312
$\lambda$-DPO 0.460 0.375 0.459 0.414 0.714 0.345
Qwen 1.8B Chat ORPO 0.377 0.278 0.417 0.347 0.661 0.007
SIMPO 0.415 0.295 0.423 0.355 0.658 0.007
DPO 0.379 0.271 0.418 0.366 0.656 0.008
$\lambda$-DPO 0.418 0.284 0.428 0.312 0.667 0.155
Table 1: Accuracy comparison across six benchmarks. The first block compares ORPO, SIMPO, DPO, and $\lambda$-DPO on LLaMA 3.2-1B Instruct. The second block compares the same four methods on Qwen 1.8B Chat. Bold values indicate the best performance within each model group.
Visualization
To further illustrate these results, we provide bar charts comparing DPO and $\lambda$-DPO performance on the two models.
Figure 2: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B Instruct.
Figure 3: Accuracy comparison between DPO and $\lambda$-DPO on Qwen 1.8B Chat.
Performance Analysis and Discussion
As shown in both Table 1 and Figures 23, $\lambda$-DPO consistently improves performance across most tasks. The method yields notable improvements on ARC, GSM8K, and MMLU, while maintaining parity on Winogrande. TruthfulQA is the only benchmark where DPO slightly outperforms $\lambda$-DPO (in the Qwen model).
These results suggest that $\lambda$-DPO effectively captures human preference signals across diverse task types. Its listwise preference modeling contributes to better alignment with human-evaluated quality metrics such as factuality, helpfulness, and fluency.
Nonetheless, for certain adversarial or logic-heavy tasks like TruthfulQA, the benefits of listwise supervision may be less reliable. Integrating $\lambda$-DPO with structured supervision—such as chain-of-thought prompting (Wei et al. 2022)—could potentially mitigate this gap.
Conclusion of Comparative Analysis
In summary, $\lambda$-DPO brings two key improvements over standard DPO. First, by adopting listwise preference supervision, it leverages full ranking information instead of pairwise comparisons, resulting in richer and lower-variance gradient signals. Second, through the introduction of the $\lambda$ vector, it supports flexible interpolation across multiple alignment objectives (e.g., helpfulness, harmlessness, factuality), enabling controllable alignment behavior and improved generalization under diverse user preferences. These two enhancements jointly contribute to the observed performance gains across benchmarks. Future work may explore hybrid strategies that combine $\lambda$-DPO with structured reasoning techniques to further boost alignment quality on complex tasks.
Ablation Experiment
To further examine the effectiveness of different preference vector settings, we conduct an ablation experiment focusing on four unimodal $\lambda$-DPO models. Each model is trained using a fixed one-hot $\lambda$ vector, targeting only one of the four annotated dimensions: helpfulness, honesty, instruction-following, and truthfulness. This setup allows us to isolate the influence of each individual preference axis and compare their respective impacts across evaluation benchmarks.
Unimodal Comparison Across Benchmarks
We evaluate the four unimodal $\lambda$-DPO variants across six standard benchmarks—mmlu, arc, hellaswag, truthfulqa, wsc273, and gsm8k. Figure 4 shows the accuracy distribution of each model across these tasks.
Figure 4: Accuracy Comparison of Unimodal $\lambda$-DPO Models with Fixed Preference Vectors.
All four models perform similarly across benchmarks. The honesty-focused model ($\lambda=[0,1,0,0]$) achieves the highest average accuracy, while the helpfulness-focused model ($\lambda=[1,0,0,0]$) excels on truthfulqa and performs well on gsm8k. Other preference models show only slight drops, with no significant performance degradation.
Preference Ranking and Vector Hypothesis
To better assess how each individual preference dimension contributes to generalization performance, we compute the average accuracy of the four unimodal $\lambda$-DPO models across six benchmark tasks:
  • $\lambda=[1,0,0,0]$: 0.4593
  • $\lambda=[0,1,0,0]$: 0.4598
  • $\lambda=[0,0,1,0]$: 0.4543
  • $\lambda=[0,0,0,1]$: 0.4548
  • $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$: 0.4612
The results indicate that all four unimodal preference dimensions contribute to model alignment, with helpfulness and honesty yielding slightly better generalization. However, the uniform interpolation $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ achieves the highest overall accuracy (0.4612), outperforming all unimodal variants.
This highlights the benefit of explicitly retaining the $\lambda$ vector for multi-objective supervision. Unlike standard DPO, which implicitly collapses preferences into a fixed form, $\lambda$-DPO provides fine-grained control over alignment trade-offs and leads to more robust generalization.
Discussion and Ablation Analysis
To further understand how individual preference dimensions influence alignment, we observe from Figure 4 that:
  • The $\lambda=[1,0,0,0]$ (helpfulness) model performs best on truthfulqa and achieves solid results on gsm8k.
  • The $\lambda=[0,1,0,0]$ (honesty) model leads slightly in overall average accuracy, especially strong on mmlu.
  • The $\lambda=[0,0,1,0]$ (instruction-following) and $\lambda=[0,0,0,1]$ (fluency) models show close and competitive performance, contrary to the notion that they are stylistically weaker.
These results indicate that optimizing for either content-driven or stylistic objectives in isolation yields comparable outcomes under our listwise framework. It further motivates the integration of multiple preference signals, especially for tasks requiring both correctness and user-aligned expressiveness.
Performance-Driven Lambda Scheduling via Polynomial Modeling
Performance-Driven Scheduler Design
We propose a performance-driven scheduler to improve multi-objective alignment by dynamically selecting preference vectors $\lambda$ based on their estimated effectiveness. Instead of fixing a single $\lambda$, the scheduler learns a distribution over $\lambda$ configurations that reflect observed performance.
The scheduler fits a polynomial regression model to past $(\lambda,\text{accuracy})$ pairs and uses it to score new candidates sampled from a Dirichlet prior over the simplex. A softmax over predicted scores then defines sampling probabilities, balancing exploitation of high-performing regions and exploration. This adaptive strategy promotes training diversity and improves robustness.
Empirical Comparison with Uniform Baseline
As shown in Figure 12, the learned scheduler achieves slightly better average accuracy (0.462 vs. 0.461) across six evaluation benchmarks when compared with a static baseline that uses the uniform configuration $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$. While the performance gain is modest, the scheduler enhances robustness and reduces overfitting by continuously adapting to observed training signals.
Importantly, this framework supports flexible preference adaptation and can be extended to personalized or task-specific settings without requiring retraining.
Figure 5: Comparison between learned $\lambda$ scheduler and uniform $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ across six alignment benchmarks. The learned scheduler achieves a slightly higher average accuracy and improves robustness by distributing preference sampling over higher-quality regions of the simplex.
Extension to LLaMA-8B
To assess the scalability of $\lambda$-DPO, we extend our experiments to the LLaMA-8B model. Across six standard benchmarks, $\lambda$-DPO yields an average relative improvement of approximately 0.3% over DPO, compared to a much larger average gain of 1.7% on the LLaMA-1B model.
Figure 6: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA-8B across six benchmarks.
As shown in Figure 6, $\lambda$-DPO achieves better alignment on the 1B model, with notably larger gains on challenging tasks such as GSM8K. This suggests that smaller models benefit more from listwise multi-preference supervision, likely because alignment signals have a stronger relative effect when model capacity is limited.
Moreover, training the 1B model only requires around 20GB of GPU memory, making it both efficient and accessible for academic labs, startups, or edge-focused applications. This low resource footprint enables rapid experimentation and deployment on commodity hardware, without sacrificing alignment quality. As such, $\lambda$-DPO is well-suited for scalable, real-world use cases where cost, speed, and controllability are critical.
Conclusion
In this paper, we propose Multi-Preference Lambda-weighted Listwise DPO ($\lambda$-DPO) to enable efficient, controllable, and fine-grained alignment of language models with diverse human preferences.
For robust and flexible alignment, we generalize pairwise supervision to listwise preference modeling, allowing the model to learn from ranked completions and capture nuanced preference structures. To support personalized and multi-objective alignment, we introduce a simplex-weighted $\lambda$ vector that dynamically balances objectives such as helpfulness, honesty, and fluency—without retraining.
To enhance generalization and robustness, we incorporate a learned scheduler that adaptively samples high-performing $\lambda$ configurations during training, improving diversity while mitigating modeling bias. Additionally, the method is designed to be resource-efficient, requiring only 20GB of GPU memory, making it practical for deployment in constrained scenarios such as on-device assistants and educational tools.
Experimental results on LLaMA 3.2-1B and Qwen-1.8B demonstrate that $\lambda$-DPO achieves consistent improvements over standard DPO, with average accuracy gains of +1.7% and +2.8%, respectively, across standard alignment benchmarks.
Acknowledgements
We thank the contributors of the original DPO framework for their foundational work, which inspired this extension. We also appreciate the open-source efforts of the Hugging Face team and the LLaMA-Factory project for enabling scalable preference-based fine-tuning.
This work was made possible through collaborative effort and self-supported compute resources by the authors: Yuhui Sun, Xiyao Wang, Zixi Li, Zhenlong Yuan and Jinman Zhao. We further thank colleagues and early readers for helpful feedback that improved the theoretical formulation and presentation.
Appendix A A Mathematical Derivations

### v6

@@ 4.3 What does $\lambda$-DPO update do?
$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=-\mathbb{E}_{x,\lambda}\bigg[$ $\sum_{i}\left(p^{\lambda}(y_{i}|x)-P_{\theta}(y_{i}|x)\right)$
$\ \ \cdot\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)\bigg]$ (10)
This gradient updates the model to increase likelihood of responses aligned with weighted human preferences, while decreasing likelihood of responses overrepresented in the reference policy. Importantly, when $\lambda$ is fixed (e.g., $\lambda=(1,0,0)$), the objective reduces to a single-objective listwise DPO loss. Due to space limitations, full derivations of the $\lambda$-DPO objective and its gradient are deferred to Appendix A.
4.4 The Pipeline
@@
In practice, we can reuse datasets where multiple preference criteria are annotated. During inference, changing $\lambda$ enables personalized alignment without modifying model parameters.
Model Method MMLU ARC HellaSwag TruthfulQA WSC273 MathQA Average
Llama 3.2-1B Instruct SIMPO 47.52 39.75 46.06 26.07 71.79 30.72 43.65
DPO 47.52 40.37 45.87 26.07 71.43 30.89 43.69
$\lambda$-DPO 48.71 45.61 46.15 29.01 75.09 32.83 46.23
Qwen 1.8B Chat SIMPO 44.29 32.27 43.11 15.91 60.81 24.56 36.83
DPO 44.30 32.72 43.17 15.91 61.54 23.38 36.84
$\lambda$-DPO 45.01 38.51 43.82 18.12 64.84 29.41 39.95
Llama 3-8B Instruct SIMPO 63.21 59.97 77.45 45.17 77.66 31.93 59.23
DPO 61.49 63.16 76.33 45.53 72.53 31.59 58.44
$\lambda$-DPO 61.03 73.25 76.44 46.63 85.35 36.87 63.26
Qwen2 7B Instruct SIMPO 68.58 68.65 62.63 45.17 76.92 39.80 60.29
DPO 68.11 69.50 62.50 45.41 75.09 42.53 60.52
$\lambda$-DPO 68.17 69.22 63.60 46.27 78.75 42.91 61.49
Table 1: Accuracy (%) comparison across six widely used benchmarks for different preference optimization methods and model backbones. Each block corresponds to a specific model, and rows compare SIMPO, DPO, and $\lambda$-DPO. Columns report performance on MMLU, ARC, HellaSwag, TruthfulQA, WSC273, and MathQA, along with the averaged score across all benchmarks. Bold values indicate the best performance within each model group.
5 Theoretical Analysis
@@
While DPO offers a principled and scalable approach to aligning language models with human feedback, its reliance on pairwise supervision imposes intrinsic limitations in expressive power, particularly in multi-objective alignment settings. In this section, we provide a theoretical analysis of this limitation and motivate our proposed extensions for better capturing and aggregating multi-dimensional human preferences.
5.1 Pairwise Limitation
@@
Given a prompt $x$ and a sampled pair of completions $y_{w}$ (preferred) and $y_{l}$ (dispreferred), Direct Preference Optimization (DPO) relies on binary preference supervision. From a listwise perspective, such supervision can be interpreted as inducing an implicit target distribution over the sampled pair:
$p^{*}(y_{i}\mid x)=\begin{cases}1&\text{if }y_{i}=y_{w},\\ 0&\text{if }y_{i}=y_{l}.\end{cases}$ (11)
This formulation is simple and computationally efficient, allowing policy optimization directly from binary human feedback.
However, treating preferences as hard assignments discards finer-grained signals, such as partial preference strength or equivalence between responses. Moreover, because gradient updates depend on a single sampled pair, the resulting optimization can exhibit high variance across iterations. Since the objective only constrains the relative preference between $y_{w}$ and $y_{l}$, other plausible or high-quality completions for the same prompt are ignored. Consequently, the expressive capacity of pairwise DPO is limited when multiple responses are comparably suitable, a limitation that becomes particularly pronounced in multi-objective alignment settings.
5.2 Pairwise Supervision as a Degenerate Listwise Objective
Equation (11) can be viewed as a degenerate form of listwise preference supervision, where the target distribution assigns all probability mass to a single completion within a sampled pair. More generally, let $\mathcal{Y}(x)=\{y_{1},\dots,y_{K}\}$ denote a set of candidate completions for prompt $x$. A listwise preference objective aims to model a target distribution $p^{*}(\cdot\mid x)$ over $\mathcal{Y}(x)$, satisfying
$\sum_{i=1}^{K}p^{*}(y_{i}\mid x)=1,\qquad p^{*}(y_{i}\mid x)\geq 0.$ (12)
Under pairwise DPO, the effective target distribution collapses to a distribution supported on only two sampled completions, with all mass assigned to the preferred one. This corresponds to an extreme point of the probability simplex in Eq. (12), with zero entropy. As a result, the optimization objective is insensitive to the relative quality of other candidates in $\mathcal{Y}(x)$, even when multiple responses are comparably suitable under certain preference trade-offs.
5.3 Incompatibility with Multi-Objective Preferences
Consider a setting with multiple preference dimensions, such as helpfulness, factuality, and safety. Let $r^{(k)}(x,y)$ denote a latent scoring function associated with the $k$-th preference criterion. Each criterion induces its own ordering over candidate responses in $\mathcal{Y}(x)$. In general, there does not exist a single scalar function $r(x,y)$ such that
$r(x,y_{i})>r(x,y_{j})$ (13)
$\quad\Longleftrightarrow\quad r^{(k)}(x,y_{i})>r^{(k)}(x,y_{j})\;\;\forall k.$
As a consequence, different preference dimensions may induce conflicting pairwise comparisons over the same response set. Reducing such multi-dimensional preference signals to a single binary comparison necessarily discards information about preference trade-offs. From this perspective, the limitation of pairwise DPO is structural: binary comparisons cannot represent preference distributions that depend on multiple, potentially competing objectives.
$\lambda$ Setting MMLU ARC HellaSwag TruthfulQA WSC273 MathQA Average
$[1,0,0,0]$ (Helpfulness) 46.36 48.91 46.22 28.03 70.70 33.53 45.63
$[0,1,0,0]$ (Honesty) 46.48 49.04 46.25 27.91 70.33 33.63 45.61
$[0,0,1,0]$ (Instruction-Following) 46.39 49.29 46.24 28.03 71.06 33.67 45.78
$[0,0,0,1]$ (Fluency) 46.26 48.84 46.22 28.15 70.33 33.37 45.53
$[0.25,0.25,0.25,0.25]$ (Uniform) 48.71 45.61 46.15 29.01 75.09 32.83 46.23
Table 2: Ablation study of different $\lambda$-weight configurations in $\lambda$-DPO, conducted on the LLaMA 3.2-1B Instruct.
6 Results
6.1 Experimental Setup
Model Selection.
We evaluate our approach on representative open-weight LLMs from both the Llama3 Touvron et al. (2023) and Qwen Qwen (2025) families. Specifically, we consider Llama 3.2-1B Instruct, Llama 3-8B Instruct, Qwen 1.8B Chat and Qwen2 7B Instruct.
Benchmark Selection.
To evaluate the effectiveness of our approach across diverse reasoning and alignment-sensitive tasks, we conduct experiments on a collection of widely used benchmarks spanning knowledge-intensive reasoning, commonsense inference, factual accuracy, and mathematical problem solving. Specifically, we include MMLU Hendrycks et al. (2021), ARC Clark and others (2018), HellaSwag Zellers et al. (2019), TruthfulQA Lin et al. (2022), WSC273 Levesque and others (2012), and MathQA Amini et al. (2019). Detailed benchmarks statistics reported in Appendix B.
Baseline Selection.
We compare our method against representative preference-based alignment baselines that are closely related to our formulation. In particular, we include DPO Rafailov et al. (2023) and SimPO Meng et al. (2024).
Training Detail.
We set the preference scaling parameter $\beta$ to 0.1, following common practice in prior work. Models are trained for three epochs with a learning rate of $5\times 10^{-6}$, using a cosine learning rate scheduler with a warmup ratio of 0.1. We use an batch size of 32, and the maximum sequence length is set to 2048 tokens. Our experiments are conducted on 4xH200 GPUs.
Dataset.
We conduct fine-tuning using the Ultrafeedback dataset Cui et al. (2023), a large-scale, instruction-following corpus designed for alignment and preference optimization. Ultrafeedback contains diverse prompts and multiple candidate responses annotated with multi-preference signals.
6.2 Main Results
We report the main experimental results in Table 1 across six widely used benchmarks. All results are reported in terms of accuracy (%), with higher values indicating better performance. Overall, $\lambda$-DPO consistently achieves the strongest average performance across all evaluated models, outperforming both SIMPO and DPO in every setting. The gains are particularly pronounced on benchmarks that are sensitive to alignment-related properties, such as TruthfulQA and WSC273, where $\lambda$-DPO yields clear and consistent improvements. This suggests that moving beyond pairwise supervision and explicitly aggregating preference signals leads to more effective alignment with nuanced human judgments.
We further observe that $\lambda$-DPO delivers robust improvements across model scales. For smaller models (e.g., Llama 3.2-1B Instruct and Qwen 1.8B Chat), $\lambda$-DPO provides substantial relative gains, indicating improved data efficiency under limited capacity. For larger models (e.g., Llama 3-8B Instruct and Qwen2-7B Instruct), $\lambda$-DPO continues to outperform pairwise baselines, demonstrating that the proposed approach scales effectively with model size. Taken together, these results validate the effectiveness and generality of $\lambda$-DPO across diverse benchmarks and model families.
A closer inspection of Table 1 shows that the performance gains of $\lambda$-DPO are most pronounced on benchmarks that are sensitive to alignment-related properties, particularly TruthfulQA and WSC273. On WSC273, $\lambda$-DPO outperforms DPO by +3.66, +3.30, +12.82, and +3.66 points for Llama 3.2-1B, Qwen 1.8B, Llama 3-8B, and Qwen2-7B, respectively, and also surpasses SIMPO by +3.30, +4.03, +7.69, and +1.83 points on the same models. Similarly, on TruthfulQA, $\lambda$-DPO improves over DPO by +2.94 (Llama 3.2-1B), +2.21 (Qwen 1.8B), +1.10 (Llama 3-8B), and +0.86 (Qwen2-7B), while exceeding SIMPO by +2.94, +2.21, +1.46, and +1.10 points, indicating more truthful and human-consistent responses across model scales. The especially large gain on WSC273 for Llama 3-8B highlights that aggregating preference signals beyond pairwise supervision can substantially improve alignment-critical commonsense judgment. In contrast, performance differences on other benchmarks, including ARC, MathQA, and the broader commonsense task HellaSwag, are generally more moderate, suggesting that the primary advantage of $\lambda$-DPO lies in better capturing nuanced human judgments without sacrificing generalization.
$\lambda$ Setting MMLU ARC HellaSwag TruthfulQA WSC273 MathQA Average
Uniform 48.71 45.61 46.15 29.01 75.09 32.83 46.23
Learned Scheduler 48.35 48.76 46.20 29.03 75.06 33.53 46.82
Table 3: Uniform vs. Learned $\lambda$, conducted on the LLaMA 3.2-1B Instruct.
6.3 Additional Analysis
Uniform vs. Unimodal Preference Weights.
We investigate the effect of different $\lambda$ configurations by comparing a uniform mixture of preference dimensions with unimodal (one-hot) alternatives, and the model is trained to optimize a single objective at a time. Table 2 reports results on LLaMA 3.2-1B Instruct. Across unimodal settings, performance is relatively similar, and no single preference dimension consistently dominates across all benchmarks. This suggests that the gains of $\lambda$-DPO are not driven by optimizing any individual objective in isolation. In contrast, the uniform mixture $\lambda=(0.25,0.25,0.25,0.25)$ achieves the best overall average performance, yielding improvements on alignment-sensitive tasks such as TruthfulQA and WSC273 while maintaining competitive accuracy on other benchmarks. Although certain unimodal configurations achieve slightly higher scores on individual tasks (e.g., instruction-following on ARC), they fail to generalize consistently across the benchmark suite. These results support our central hypothesis that explicitly aggregating multi-dimensional preference signals leads to more robust and balanced alignment than optimizing a single preference dimension alone.
Uniform vs. Learned $\lambda$.
Static uniform $\lambda$ configurations face two key limitations. First, different benchmarks exhibit different sensitivities to preference dimensions, whereas a uniform $\lambda$ implicitly assumes that all objectives are equally important. This mismatch prevents the model from adapting to benchmark-specific preference structure and can lead to suboptimal performance. Second, although we fit a polynomial performance model $f(\lambda)$ to approximate downstream accuracy over the simplex, this approximation is inherently imperfect: $\lambda$ values predicted as weak by $f$ may still yield strong true performance. Thus, selecting only the maximizer of $f$ risks collapsing to a poorly explored region of the preference space.
Instead of committing to a single $\lambda$, we construct a stochastic scheduler that samples preference weights in a performance-aware manner. Specifically, we define an idealized continuous sampling distribution over the simplex
$p(\lambda)\;\propto\;\exp(\tau f(\lambda)),\qquad\lambda\in\Delta_{4},$
where $\tau>0$ controls the exploration–exploitation balance. Since this density cannot be computed in closed form, we discretize it by sampling candidate $\{\lambda^{(j)}\}_{j=1}^{k}$ from a Dirichlet prior and applying a softmax:
$p(\lambda^{(j)})=\frac{\exp(\tau f(\lambda^{(j)}))}{\sum_{l}\exp(\tau f(\lambda^{(l)}))}.$
This produces a learned $\lambda$ scheduler that prioritizes promising trade-offs while still exploring alternative preference allocations, explicitly guarding against misspecification of $f$.
The result is presented in Table 3, with the full formulation and implementation details of the learned $\lambda$ scheduler provided in Appendix C. Empirically, this design yields clear gains. Compared to the uniform mixture $\lambda=(0.25,0.25,0.25,0.25)$, which achieves an average score of $46.23$, the learned scheduler improves the overall average to $46.82$. It further delivers meaningful benchmark-specific improvements, such as ARC ($45.61\rightarrow 48.76$) and MathQA ($32.83\rightarrow 33.53$), while maintaining strong alignment-sensitive performance on TruthfulQA and WSC273. These results demonstrate that performance-driven stochastic $\lambda$ adaptation is more robust than any static weighting scheme.
Benchmark-Specific Preference Sensitivity.
Using the results reported in Table 2, we further analyze how different benchmarks respond to individual preference dimensions. Rather than focusing on overall performance aggregation, this analysis adopts a benchmark-centric perspective and examines which preference objectives are most aligned with specific evaluation tasks.
We observe that no single preference dimension dominates across all benchmarks. For example, ARC exhibits its strongest performance under the instruction-following–focused configuration, suggesting that precise adherence to task instructions plays a critical role in this benchmark. In contrast, TruthfulQA benefits more consistently from balanced or honesty-oriented preference mixtures, reflecting its sensitivity to factual correctness and avoidance of hallucination. Similarly, WSC273 shows improved performance under configurations that emphasize balanced reasoning and fluency, rather than any single isolated objective. These observations highlight that different benchmarks implicitly encode distinct preference structures
7 Conclusions
In this work, we revisit preference optimization for language model alignment from a multi-dimensional perspective. We identify that existing DPO-based methods implicitly assume a single fixed preference objective, which limits their expressive power when human judgments are structured, heterogeneous, and sometimes conflicting across dimensions. To address this limitation, we propose $\lambda$-weighted Listwise Direct Preference Optimization ($\lambda$-DPO), a unified framework that combines listwise supervision with flexible preference aggregation for multi-dimensional alignment. Extensive experiments across multiple model families and six widely used benchmarks demonstrate that $\lambda$-DPO consistently outperforms strong DPO-based baselines.
8 Limitations
While our method improves the flexibility and expressiveness of preference optimization, this work still has several limitations that warrant further investigation. Our evaluation primarily relies on accuracy over standard benchmarks as a proxy for alignment quality. Although these benchmarks capture important aspects such as reasoning ability, factual correctness, and instruction-following, they do not fully reflect the complexity of real-world deployment scenarios, including long-horizon interactions, dynamically changing user preferences, and safety-critical behaviors.
Appendix A Mathematical Derivations